OpenStudy (anonymous):

If f(x) = x^2 +5x, find lim delta x f (x + delta x ) - f (x) / delta x

OpenStudy (anonymous):

\[\lim_{\Delta x\to0}\frac{f(x+\Delta x)-f(x)}{\Delta x}\] You left it out, but I'm assuming \(\Delta x\) is approaching zero, since this looks like you're supposed to find the derivative of the given \(f(x)\) using the limit definition. Also, for the sake of brevity, I'm going to denote \(h=\Delta x\). \[\begin{align*}\lim_{h\to0}\frac{f(x+h)-f(x)}{h}&=\lim_{h\to0}\frac{\bigg((x+h)^2+5(x+h)\bigg)-\bigg(x^2+5x\bigg)}{h}\\ &=\lim_{h\to0}\frac{\bigg(x^2+2xh+h^2+5x+5h\bigg)-\bigg(x^2+5x\bigg)}{h}\\ &=\lim_{h\to0}\frac{2xh+h^2+5h}{h}\\ &=\lim_{h\to0}(2x+h+5)\\ &=\cdots \end{align*}\]